Conditions: 0cm of new snow in the past 24 hours, 40cm in the past 7 days, total snowpack heights averaging between 170-200cm in most riding areas.
- Hartley: Groomed from the staging area to the East Fork of the Sulphur last night, parking lot plowed, a significant number of trees blocking trail were removed
- Morrissey: Morrissey Trail, Doug Dean, McEvoy Trail, South Pipeline Trail, Summit Trail, and Notch trail groomed Jan 14th
- Corbin: Last groomed Jan 8nd (Friday each week)
- Coal Creek: Remains CLOSED to public parking and sledding for timber harvesting operations
NOTES: COVID signage posted in cabins is a requirement of our continued operation. If individuals continue to remove signage cabins will be locked and closed to the public.